A shipping box measures 14 1/2 inches long by 9 1/8 inches wide by 8 11/16 inches high. What is the volume of the box in cubic feet? Round to the nearest hundredth of a cubic foot.: A shipping box measures 14 1/2 inches long by 9 1/8 inches wide by 8 11/16 inches high. What is the volume of the box in cubic feet? Round to the nearest hundredth of a cubic foot.
Answer by Edwin McCravy(2087)

A shipping box measures 14 1/2 inches long by 9 1/8 inches wide by 8 11/16 inches high. What is the volume of the box in cubic feet? Round to the nearest hundredth of a cubic foot.

Change 141/2in to the decimal 14.5in
Change 91/8in to the decimal 9.125in
Change 811/16in to the decimal 8.6875in

V=LWH

V=(14.5in)(9.125in)(8.6875in)

V=(14.5)(9.125)(8.6875)(in^3)

V=1149.464844(in)^3

Now put this over 1

V=1149.464844(in)^3/1

Now we make a unit fraction, using
inches and feet that is a fraction 
that equals 1, that has equal
numerator and denominator, and which
will cancel away the inches. We know
that 12in=1ft. Therefore such a
unit fraction is (1(ft))/(12(in)) 

But we need to cancel away (in)^3,
so we cube the unit fraction and get
(1^3(ft)^3)/(12^3(in)^3) or
(1(ft)^3)/(1728(in)^3)

Now we multiply by that unit fraction:

V=(1149.464844(in)^3/1)((1(ft)^3)/(1728(in)^3))

Now we can cancel the (in)^3

V=(1149.464844cross((in)^3)/1)((1(ft)^3)/(1728cross((in)^3)))

and all we have left is

V=(1149.464844(ft)^3)/(1728)

Divide that out and get:

V=0.6651995624ft^3

or to the nearest hundredth of a ft^3

V=0.67ft^3

Edwin